# Evidence of partonic collectivity in ultra-relativistic heavy-ion collisions with NCQ scaling of radial flow

### v1

@@ 3.2 Observable
{tikzpicture}
@@
[…]
Figure 2: Cartoon illustrating the $\eta$ regions used in the analysis. Regions A (red) and B (green) are separated by a gap $\eta_{\rm gap}$. The observable $F_{\rm A}^{p_{\rm T}}$ is correlated with the transverse momentum $p_{\rm T}]_{\rm B}$ measured in the opposite region.
The term $\sigma_{[p_{\rm T}]}$ in Eqn. 1 can be expressed as follows,
Here, $\langle..\rangle$ represents event-average and $\sigma_{[p_{\rm T}]}$ is the standard deviation of mean-$p_{\rm T}$ fluctuations.
Figure 4: (Color online) Comparison of $v_{0}(p_{\rm T})$ between PYTHIA8/Angantyr predictions and ALICE measurements [10] for identified hadrons in Pb+Pb collisions at = $\sqrt{s_{\rm NN}}$ = 5.02 TeV, presented for 10-20% (left panel) and 30-40% (right panel) centrality classes. Model calculations are represented by solid curves with shaded bands denoting bootstrap uncertainties, while experimental data are shown as filled markers.
@@
Figure 7: (Color online) NCQ-scaled $v_{0}(p_{\rm T})$ of identified hadrons as a function of NCQ-scaled $p_{\rm T}$ (top panel) and as a function of NCQ-scaled ($m_{\rm T}-m_{\rm 0}$) (bottom panel) in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. This plot is compiled using the experimental data of ALICE [10].
4 Results and Discussion
4.1 Estimation of $v_{0}(p_{\rm T})$ at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
In Fig. 1, we display the radial flow observable, $v_{0}(p_{\rm T})$ for identified hadrons ($\pi^{\pm}$, $K^{\pm}$, and $p+\overline{p}$) in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, obtained from AMPT String Melting simulations for central (0-20%) and semi-central (20-40%) class. The observable is calculated using the subevent method with $\eta_{\rm gap}$ = 0.4 to minimize the short-range non-flow correlations. A clear mass ordering is observed at lower-$p_{\rm T}$ ($p_{\rm T}\leq 2$ GeV/c), while a characteristic baryon–meson crossover emerges at intermediate $p_{\rm T}$ ($2\leq p_{\rm T}\leq 3$ GeV/c). These features are consistent with expectations from collective radial expansion of the produced medium. The trends align with measurements at $\sqrt{s_{\rm NN}}=5.02$ TeV by ALICE [10] and with hydrodynamic model calculations [17].
[…]
4.2 Estimation of $v_{0}(p_{\rm T})$ at $\sqrt{s_{\rm NN}}=5.02$ TeV
This analysis is extended to Pb+Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV by comparing the radial flow observable $v_{0}(p_{\rm T})$ for identified charged hadrons from PYTHIA8/Angantyr simulations with the corresponding ALICE data [10] for 10–20% (central) and 30–40% (semi-central) classes, as shown in Fig. 4. As reported by the ALICE collaboration [10], the experimental results display clear signatures of collective expansion, characterized by a pronounced mass ordering at low transverse momentum ($p_{\rm T}\leq 3~\text{GeV}/\textit{c}$) and a distinct meson-baryon crossing at intermediate transverse momentum ($p_{\rm T}\geq 3~\text{GeV}/\textit{c}$). In contrast, the $v_{0}(p_{\rm T})$ values predicted by PYTHIA8/Angantyr remain close to zero over the entire $p_{\rm T}$ range, exhibiting neither mass ordering nor meson-baryon crossing. This pronounced discrepancy reflects the absence of a hydrodynamic evolution stage in the model, which prevents the buildup of sustained pressure gradients required to generate collective radial flow. The inability of the model to reproduce the observed experimental features provides further evidence that the measured $v_{0}(p_{\rm T})$ patterns arise from genuine final-state collective dynamics, rather than from initial-state effects or string-fragmentation mechanisms.
[…]
4.3 NCQ scaling of radial flow observable $v_{0}(p_{\rm T})$
4.3.1 NCQ scaling at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$

### v2

@@ III.2 Observable
The term $\sigma_{[p_{\rm T}]}$ in the denominator of Eq. (1) is the standard deviation of mean-$p_{\rm T}$ and can be calculated as follows,
The integrated observable $v_{0}$ quantifies event-by-event fluctuations of the mean transverse momentum $p_{\rm T}]$. Following Ref. [2], and based on the fluctuation formalism of Ref. [29], it is defined as the relative root-mean-square of the fluctuations of the event-wise mean transverse momentum evaluated within a chosen reference transverse-momentum range $R$,
$v_{0}\equiv\frac{\sqrt{\langle(\delta[p_{\rm T}])^{2}\rangle_{R}}}{\langle[p_{\rm T}]\rangle_{R}},$ (5)
where $\delta[p_{\rm T}]=[p_{\rm T}]-\langle[p_{\rm T}]\rangle_{R}$, and $\langle\cdots\rangle_{R}$ denotes averaging over events with particles restricted to the reference interval $R$.
Under the assumption of collective dynamics, fluctuations of the single-particle spectrum $\delta n(p_{\rm T})$ are linearly correlated with fluctuations of $p_{\rm T}]$, leading to an approximate factorization of the normalized covariance. This yields a sum rule relating the integrated observable to its transverse-momentum-differential counterpart $v_{0}(p_{\rm T})$,
$v_{0}\sim\frac{\int_{p_{\rm T}\in R}(p_{\rm T}-\langle[p_{\rm T}]\rangle_{R})\langle n(p_{\rm T})\rangle v_{0}(p_{\rm T})\,dp_{\rm T}}{\int_{p_{\rm T}\in R}p_{\rm T}\langle n(p_{\rm T})\rangle\,dp_{\rm T}},$ (6)
where $\langle n(p_{\rm T})\rangle$ denotes the event-averaged single-particle spectrum.
Figure 4: (Color online) Centrality dependence of $v_{0}(p_{\rm T})$ (left), $v_{0}(p_{\rm T})/v_{0}$ (center) and $v_{0}(p_{\rm T})\sqrt{\rm N_{ch}}$ (right), presented as a function of $p_{\rm T}$ for Au+Au simulations using AMPT-SM at $\sqrt{s_{\rm NN}}=200$ GeV. A reference-$p_{\rm T}$ range of 0.2-1.5 GeV is selected for this particular analysis.
IV Results and Discussion
IV.1 Transverse momentum dependence of $v_{0}(p_{\rm T})$
Fig. 3 shows the $p_{\rm T}$ dependence of the radial flow observable $v_{0}(p_{\rm T})$ for 0–5% (most central), 30–40% (mid-central), and 70–80% (peripheral) centrality classes in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, computed using the AMPT-SM model. The calculations employ a pseudorapidity gap $\eta_{\rm gap}=0.4$ within $|\eta|\leq 1.0$ to suppress short-range non-flow correlations. At low-$p_{\rm T}$ ($\lesssim 0.5$ GeV/$c$), $v_{0}(p_{\rm T})$ is negative across all centrality intervals, consistent with the anti-correlation between event-by-event (EbyE) mean-$p_{\rm T}$ fluctuations and the differential particle yield at a given $p_{\rm T}$ [29, 26]. This behavior reflects that events with larger-than-average mean transverse momentum ($p_{\rm T}]>\langle[p_{\rm T}]\rangle$), corresponding to stronger radial expansion, produce flatter $p_{\rm T}$ spectra with enhanced yields at high-$p_{\rm T}$ and suppressed yields at low-$p_{\rm T}$, while events with weaker radial expansion exhibit the opposite trend [26]. For $p_{\rm T}>0.55$ GeV/$c$, $v_{0}(p_{\rm T})$ turns positive and rises non-linearly with $p_{\rm T}$. The magnitude of $v_{0}(p_{\rm T})$ is significantly larger for the 70–80% centrality class than for 0–5% and 30–40%, consistent with the expectation that peripheral collisions exhibit larger EbyE geometric fluctuations, which drive stronger radial flow fluctuations [16, 1]. Crucially, the three $p_{\rm T}^{\rm ref}$ ranges yield mutually consistent results for $p_{\rm T}\lesssim 1.6$ GeV/$c$, demonstrating that $v_{0}(p_{\rm T})$ factorizes from the choice of reference range in this regime (a direct signature of collective radial flow) while deviations emerging at higher-$p_{\rm T}$ indicate the onset of non-flow contributions [2, 3]. One can see that $v_{0}(p_{\rm T})$ magnitude for $p_{\rm T}^{\rm ref}:0.2-1.5$ GeV dips at higher-$p_{\rm T}$, indicating that it is least affected by the non-flow correlations.
IV.2 Centrality dependence of $v_{0}(p_{\rm T})$
Fig. 4 (left) presents the $p_{\rm T}$ dependence of $v_{0}(p_{\rm T})$ for inclusive charged hadrons in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, obtained from AMPT-SM simulations across nine centrality classes viz. 0-5%, 5-10%, 10-20%, 20-30%, 30-40%, 40-50%, 50-60%, 60-70%, and 70-80%. A clear centrality dependence is observed over the entire $p_{\rm T}$ range. The most central events (0–5%) exhibit the smallest values of $v_{0}(p_{\rm T})$, whereas the most peripheral class (70–80%) shows the largest magnitude, particularly at higher-$p_{\rm T}$. The increase of $v_{0}(p_{\rm T})$ with $p_{\rm T}$ is modest for central collisions, whereas a more pronounced rise is observed in peripheral events. This systematic evolution can be understood in terms of the interplay between collision geometry and collective expansion dynamics. Although central collisions generate larger and longer-lived systems with stronger average radial flow, the relative magnitude of flow-induced fluctuations, as quantified by $v_{0}(p_{\rm T})$, is reduced due to the larger participant multiplicity, which effectively averages out event-by-event fluctuations.
Fig. 4 (center) shows the $p_{\rm T}$ dependence of the scaled radial flow observable $v_{0}(p_{\rm T})/v_{0}$ for different centrality classes, where $v_{0}$ denotes the corresponding $p_{\rm T}$-integrated radial flow magnitude (See Eq. (5)). The normalized distributions $v_{0}(p_{\rm T})/v_{0}$ exhibit a clear centrality-independent scaling behavior up to $p_{\rm T}\lesssim 2$ GeV/$c$. This scaling suggests that, in this momentum region, the correlations are predominantly governed by bulk collective dynamics within the AMPT-SM framework. A qualitatively similar scaling trend has been reported by the ATLAS Collaboration in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [2], indicating that this feature persists across collision energies. At higher transverse momentum, $p_{\rm T}\gtrsim 2$ GeV/$c$, systematic deviations from the low-$p_{\rm T}$ scaling trend become apparent, particularly in peripheral collisions where $v_{0}(p_{\rm T})/v_{0}$ continues to increase rather than following the universal behavior observed at lower-$p_{\rm T}$. These deviations are consistent with a reduced dominance of collective expansion and an increasing relative contribution from non-collective particle production mechanisms, such as hard partonic scatterings and jet fragmentation.
To further probe the origin of radial flow fluctuations, we examine $v_{0}(p_{\rm T})\sqrt{\rm N_{\rm ch}}$ as a function of $p_{\rm T}$ for different centralities in Fig. 4 (right). An approximate collapse of the curves is observed at low-$p_{\rm T}$ (up to $\sim 1.2$ GeV/$c$), consistent with a $1/\sqrt{\rm N_{\rm ch}}$ scaling in this momentum region. A similar trend has been reported in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV and interpreted within an effective independent source framework [2]. In the present AMPT-SM calculations, which incorporate partonic scatterings followed by quark coalescence and hadronic rescattering, the same scaling behavior persists despite the presence of strong interactions in the partonic phase. The persistence of the scaling in this interacting system suggests that EbyE fluctuations of the collective radial expansion are progressively diluted with increasing multiplicity. At higher-$p_{\rm T}$, the scaling gradually breaks down, reflecting the increasing contribution of harder partonic processes that are less governed by bulk collective dynamics.
Figure 5: (Color online) Centrality dependence of integrated radial flow measure, $v_{0}$ (left panel) and $v_{0}/v_{0}^{0-5\%}$ (right panel) for inclusive charged hadrons at three different $p_{\rm T}^{\rm ref}$ ranges for Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV simulated using AMPT-SM framework.
Figure 6: (Color online) Pseudorapidity gap dependence of inclusive $v_{0}(p_{\rm T})$ within three centrality intervals: 0-10% (most central), 30-40% (semi-central) and 70-80% (peripheral) for Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV simulated using the AMPT-SM framework.
Figure 7: (Color online) Dependence of $v_{0}(p_{\rm T})$ on the choice of centrality estimators for most central, mid-central and peripheral Au+Au collisions generated using AMPT-SM at $\sqrt{s_{\rm NN}}=200$ GeV.
IV.3 Centrality dependence of integrated radial flow fluctuations
The top left panel of Fig. 5 presents the centrality evolution of the integrated radial flow observable $v_{0}$ (calculated using Eq. (5)) in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV for three reference transverse momentum ranges: $p_{\rm T}^{\rm ref}=0.2-1.5~\rm GeV,0.2-2.0~\rm GeV$ and $0.5-2.0~\rm GeV$. The results reveal a systematic increase in $v_{0}$ from central to peripheral collisions across all $p_{\rm T}^{\rm ref}$ selections, consistent with the expected behavior of event-by-event (EbE) radial flow fluctuations in heavy-ion collisions [2, 12]. The magnitude of $v_{0}$ exhibits a clear dependence on the chosen reference momentum window, with the widest range ($0.2-2.0~\rm GeV$) yielding the largest values. This behavior follows directly from the sum rule relating $v_{0}$ to $v_{0}(p_{\rm T})$, where broader integration ranges encompass larger portions of the momentum spectrum affected by radial flow fluctuations [2]. The range ($0.5-2.0~\rm GeV$) produces correspondingly smaller $v_{0}$ values, reflecting the reduced contribution from low-momentum particles where radial flow effects are most pronounced.
The top right panel of Fig. 5 shows $v_{0}$ normalized to its value in the 0–5% most central collisions, $v_{0}/v_{0}^{0-5\%}$. The three $p_{\rm T}^{\rm ref}$ selections exhibit similar normalized trends, with the corresponding curves overlapping within systematic uncertainties. This approximate scaling indicates that although the absolute magnitude of radial flow fluctuations depends on the chosen momentum interval, the centrality dependence follows a common underlying pattern that is largely independent of the specific $p_{\rm T}^{\rm ref}$ selection. Within the string melting configuration of AMPT, this scaling can be attributed to event-by-event fluctuations in the effective transverse expansion velocity generated during the partonic cascade stage. Fluctuations in the initial parton density and subsequent scattering history coherently modify the transverse momentum spectra during the partonic cascade stage. Hence, $v_{0}$ captures event-by-event fluctuations of the global spectral shape, consistent with a collective origin of radial flow fluctuations generated in the partonic evolution. The present findings are consistent with the results reported for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV at the LHC by the ATLAS Collaboration [2], where a similar scaling behavior of $v_{0}$ was observed, supporting its interpretation as a measure of collective radial flow fluctuations.
The same analysis is presented in the bottom panels of Fig. 5, where $\rm N_{ch}$ is used as the centrality variable. Since $\rm N_{ch}$ is inversely correlated with centrality, $v_{0}$ exhibits a monotonic decrease with increasing multiplicity, consistent with the trends observed in the centrality-based representation. The dependence on the chosen $p_{\rm T}^{\rm ref}$ interval remains unchanged, with the widest momentum range yielding the largest magnitude of $v_{0}$. The bottom right panel shows $v_{0}$ normalized to its value at $\rm N_{ch}^{5\%}$ as a function of $\rm N_{ch}/N_{ch}^{5\%}$. The three $p_{\rm T}^{\rm ref}$ selections again display approximate scaling within uncertainties, demonstrating that the multiplicity-based representation preserves the same underlying fluctuation pattern. Quantitatively, the multiplicity dependence is well described by a power-law parametrization, $v_{0}=AN_{\rm ch}^{b}$, for all $p_{\rm T}^{\rm ref}$ intervals. The extracted exponents, summarized in Tab. 2, lie in the range 0.42-0.42 to 0.45-0.45 with only weak momentum dependence. For comparison, the corresponding measurements reported by the ATLAS Collaboration were fitted with the same functional form (See Fig. 15), yielding exponents close to $b\approx-0.40$ to 0.42-0.42. In both cases, the magnitude of $b$ is consistent with an approximate inverse square-root dependence on $\rm N_{\rm ch}$, indicating that relative radial flow fluctuations decrease with increasing multiplicity. The overall agreement in scaling behavior between model and data supports a common multiplicity-driven origin of radial flow fluctuations, consistent with the collective nature of the underlying transverse expansion.
Table 2: Power-law fit parameters for the $v_{0}$ dependence in AMPT-SM Au+Au collisions (top) and ATLAS Pb+Pb data [2] (bottom), obtained using $v_{0}=A\,N_{\rm ch}^{\,b}$.
System ($\sqrt{s_{\rm NN}}$) $p_{\rm T}^{\rm ref}$ [GeV] AA $b$
0.20.21.51.5 $0.130\pm 0.001$ $0.424\pm 0.002$
Au+Au ($200$ GeV) 0.20.22.02.0 $0.164\pm 0.001$ $0.438\pm 0.001$
0.50.52.02.0 $0.080\pm 0.002$ $0.452\pm 0.005$
0.50.52.02.0 $0.185\pm 0.011$ $0.421\pm 0.008$
Pb+Pb ($5.02$ TeV) 0.50.55.05.0 $0.316\pm 0.017$ $0.411\pm 0.007$
1.01.05.05.0 $0.204\pm 0.011$ $0.408\pm 0.007$
IV.4 $\eta_{\rm gap}$ dependence of $v_{0}(p_{\rm T})$
To isolate the contributions from short-range correlations, we investigate the pseudorapidity gap dependence of $v_{0}(p_{\rm T})$ in Fig. 6 for three centrality classes (0–10%, 30–40%, and 70–80%). The observable exhibits the characteristic sign change, being negative at low-$p_{\rm T}$ (<0.5<0.5 GeV/$c$) and positive at higher-$p_{\rm T}$ (>0.5>0.5 GeV/$c$). The weak dependence on $\eta_{\rm gap}$, particularly in central and mid-central collisions, indicates that $v_{0}(p_{\rm T})$ is predominantly driven by long-range collective correlations rather than short-range nonflow effects. In peripheral collisions (70–80%), a mild $\eta_{\rm gap}$ dependence emerges at higher-$p_{\rm T}$, suggesting a modest contribution from jet-like correlations where the system size and collective effects are reduced. These observations support the interpretation that radial flow in Au+Au collisions at RHIC energies exhibits collective behavior consistent with recent measurements in Pb+Pb collisions at the LHC [3, 2].
IV.5 Effect of centrality estimator on $v_{0}(p_{\rm T})$
In this section, we compare two commonly used centrality estimators: midrapidity charged-particle multiplicity ($|\eta|\leq 0.5$) and the impact parameter $b$. The impact parameter directly characterizes the geometric overlap of the colliding nuclei, whereas the multiplicity-based estimator mimics the experimental procedure in which centrality is determined from particle yields measured within a given detector acceptance. Fig. 7 shows $v_{0}(p_{\rm T})$ as a function of $p_{\rm T}$ for three representative centrality classes (0–10%, 30–40%, and 70–80%) obtained using both estimators. We observe that at low-$p_{\rm T}$, $v_{0}(p_{\rm T})$ is nearly identical for both centrality estimators across all centrality classes. At higher-$p_{\rm T}$, small but systematic differences emerge, with the impact parameter based selection yielding slightly smaller values of $v_{0}(p_{\rm T})$. This behavior can also be understood in terms of residual volume fluctuations associated with multiplicity-based centrality selection. Within a given multiplicity interval, event-by-event variations in entropy deposition lead to fluctuations in the effective system size, which couple to variations of the spectral shape. Such volume-induced fluctuations can modestly enhance the extracted $v_{0}(p_{\rm T})$, particularly at high-$p_{\rm T}$ where relative spectral modifications induced by variations in collective transverse expansion become more pronounced. In contrast, selecting events by the impact parameter constrains the collision geometry more directly and reduces the influence of residual volume fluctuations.
Figure 8: (Color online) Transverse momentum ($p_{\rm T}$) dependence of $v_{0}(p_{\rm T})$ for identified hadrons in Au+Au collisions generated with the String Melting version of AMPT model, shown for 0-20% (most central) and 20-40% (semi-central) centrality selections. Shaded bands indicate bootstrap-estimated statistical uncertainties.
Figure 10: (Color online) Comparison of $v_{0}(p_{\rm T})$ between PYTHIA8/Angantyr predictions and ALICE measurements [3] for identified hadrons in Pb+Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV, presented for 10-20%, 30-40% and 60-70% centrality intervals. Model calculations are represented by solid curves with shaded bands denoting bootstrap uncertainties, while experimental data are shown as filled markers with uncertainties shown in vertical bars (statistical) and hollow boxes (systematic).
IV.6 $v_{0}(p_{\rm T})$ for identified charged hadrons at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
In Fig. 8, we display the radial flow observable, $v_{0}(p_{\rm T})$ for identified hadrons ($\pi^{\pm}$, $K^{\pm}$, and $p+\overline{p}$) in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, obtained from AMPT-SM simulations for central (0-20%) and mid-central (20-40%) class. The observable is calculated using the subevent method with $\eta_{\rm gap}$ = 0.4 to minimize the short-range non-flow correlations. A clear mass ordering is observed at lower-$p_{\rm T}$ ($p_{\rm T}\leq 2$ GeV/c), while a characteristic baryon–meson crossover emerges at intermediate $p_{\rm T}$ ($2\leq p_{\rm T}\leq 3$ GeV/c). These features are consistent with expectations from collective radial expansion of the produced medium. The trends align with measurements at $\sqrt{s_{\rm NN}}=5.02$ TeV by ALICE [3] and with hydrodynamic model calculations [18].
@@ IV.6 $v_{0}(p_{\rm T})$ for identified charged hadrons at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
IV.7 Estimation of $v_{0}(p_{\rm T})$ at $\sqrt{s_{\rm NN}}=5.02$ TeV
This analysis is further extended to Pb+Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV by comparing the radial flow observable $v_{0}(p_{\rm T})$ for identified charged hadrons from PYTHIA8/Angantyr simulations with the corresponding ALICE data [3] for 10–20% (central), 30–40% (semi-central) and 60-70% (peripheral) classes, as shown in Fig. 10. As reported by the ALICE collaboration [3], the experimental results display clear signatures of collective expansion, characterized by a pronounced mass ordering at low-$p_{\rm T}$ ($p_{\rm T}\leq 3~\text{GeV}/\textit{c}$) and a distinct meson-baryon crossing at intermediate-$p_{\rm T}$ ($p_{\rm T}\geq 3~\text{GeV}/\textit{c}$). In contrast, the $v_{0}(p_{\rm T})$ values predicted by PYTHIA8/Angantyr remain close to zero over the entire $p_{\rm T}$ range, exhibiting neither mass ordering nor meson-baryon crossing. This pronounced discrepancy reflects the absence of a hydrodynamic evolution stage in the model, which prevents the buildup of sustained pressure gradients required to generate collective radial flow. The inability of the model to reproduce the observed experimental features provides further evidence that the measured $v_{0}(p_{\rm T})$ patterns arise from genuine final-state collective dynamics, rather than from initial-state effects or string-fragmentation mechanisms.
@@
Figure 13: (Color online) NCQ-scaled $v_{0}(p_{\rm T})$ of identified hadrons as a function of NCQ-scaled $p_{\rm T}$ (top panel) and as a function of NCQ-scaled ($m_{\rm T}-m_{\rm 0}$) (bottom panel) in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. These plots are compiled using the experimental data of ALICE [3]. Vertical bars represent statistical uncertainties.
IV.8 NCQ scaling of radial flow observable $v_{0}(p_{\rm T})$
IV.8.1 NCQ scaling at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$